OpenStudy (anonymous):

How would you solve a problem like this: Evaluate the logarithmic expression. Do not round the expression until the final answer. When necessary, round to the nearest hundredth. log6 21

OpenStudy (ipwnbunnies):

I'm assuming that's base 6. So what this is saying is.... 6^x = 21. We're solving for x, and it's not too easy to find at first glance. What you can do is use a calculator and do this: log(21)/log(6)

OpenStudy (anonymous):

Thank you!! My brain is fried from 3 hours of constantly changing math problems and I couldn't think of this. haha

OpenStudy (anonymous):

Express in exponential form. log 3 a = n ? Help with this one?

OpenStudy (ipwnbunnies):

log base 3 (a) = n. I'll put it in words to see if you can write it. The base, 3, raised to some power, n, will equal a.

OpenStudy (anonymous):

Thanks! Okay, one more for you: log 10 -2 =

OpenStudy (ipwnbunnies):

log base 10 of -2 = ?. K, 10^x = -2. This does not exist b/c x can be anything, you will not get a negative answer.

OpenStudy (ipwnbunnies):

The logarithm of a negative number does not exist.

OpenStudy (anonymous):

So, would the answer be 2 or 0? Those are my choices.

OpenStudy (ipwnbunnies):

Neither lol. Did you write out the problem correctly?

OpenStudy (anonymous):

I am taking a quiz on an online class course and the answers I am given are -2, 2 and 0. But you said it wasn't going to be a negative! haha

OpenStudy (ipwnbunnies):

I said you can't take the logarithm of a negative number, so there can't be an answer. D:

OpenStudy (ipwnbunnies):

Are you certain you wrote it correctly, it's not log 10 = -2?

OpenStudy (anonymous):

nope, I copied and pasted the exact problem. I guessed, and I made a 58.6 on the quiz....haha Now I have another test over this to take....

OpenStudy (ipwnbunnies):

Lol, I'm not sure then. That was a dumb question.

OpenStudy (anonymous):

Would you like to help me on this test as well? :P

OpenStudy (ipwnbunnies):

If it's easy. Idk why I'm on here, I don't want to do my Calculus 3 hmwk. D:

OpenStudy (ipwnbunnies):

I'm not helping if I have to write stuff lol

OpenStudy (anonymous):

It will be pretty much the same type questions. haha.

OpenStudy (ipwnbunnies):

Ok

OpenStudy (anonymous):

Some of them I will know, because, I'm not completely retarded, I swear! haha

OpenStudy (ipwnbunnies):

lol :p

OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (ipwnbunnies):

Is standard form y = mx + b? Or you have x and y on both sides....

OpenStudy (anonymous):

I think standard is y=mx+b

OpenStudy (ipwnbunnies):

We need to find the slope created by these two points. Eq for slope: \[m = \frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

so, -2 and -3

OpenStudy (ipwnbunnies):

Then we use point-slope equation to find the eq. for the line: (y-y1) = m(x-x1) I think you can either use (1,5) or (-2,3)

OpenStudy (ipwnbunnies):

Then divide them to get the slope of the line through those points. So m = 2/3

OpenStudy (anonymous):

Hmm, that is the only question I needed actually, the rest is just graphing! thanks :D

OpenStudy (ipwnbunnies):

Sweet, good luck.

OpenStudy (anonymous):

Hopefully I don't fail this one too! xD Go do your Calculus 3 work! :P